# From Zipf‘s Law to Neural Scaling through Heaps‘ Law and Hilberg‘s Hypothesis

### v1

@@ 1 Introduction
In the present paper, we aim to provide an actually simple but systematic derivation of this chain of implications that extends earlier results in information theory and quantitative linguistics [23, 25, 26, 28]. We will prove formal theorems about as general stochastic processes as possible rather than experiment with particular empirical data. Our goal is to consolidate knowledge of several scientific disciplines by means of mathematical deduction.
The conceptual trajectory is as follows. Beginning with Zipf‘s law for word frequencies, we derive Heaps‘ law for vocabulary growth. From Heaps‘ law we extract a formal version of Hilberg‘s hypothesis on the sublinear growth of block entropy. Finally, we show that Hilberg‘s hypothesis leads to the neural scaling that ties model performance to training data size, model size, and amount of training compute. In parallel, we will discuss an example of Santa Fe processes—toy stochastic sources that exhibit these statistical laws. Each link in this chain introduces its own assumptions and modeling choices. By isolating these steps, we hope to illuminate where the derivations rely on idealizations and how they might be strengthened.
@@
Notation $(X_{t})_{t\in\mathbb{Z}}$ denotes an integer-time countable-alphabet stochastic process, $X_{j}^{k}:=(X_{j},X_{j+1},\ldots,X_{k})$ is a string of random tokens, and $\mathcal{X}_{j}^{k}:=\left\{X_{j},X_{j+1},\ldots,X_{k}\right\}$ is a random set of types. We adopt that $X_{1}^{0}$ is the empty string and $\mathcal{X}_{1}^{0}$ is the empty set. Notation $H(X):=\operatorname{\mathbf{E}}(-\log P(X))$ stands for the Shannon entropy of a random variable and $V(\mathcal{X}):=\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{X}$ denotes the expected cardinality of a random set. The conditional entropy is H(X|Q):=H(X,Q)H(Q)H(X|Q):=H(X,Q)-H(Q), the mutual information is I(X;Y):=H(X)H(X|Y)I(X;Y):=H(X)-H(X|Y), and the conditional mutual information is I(X;Y|Q):=H(X|Q)H(X|Y,Q)I(X;Y|Q):=H(X|Q)-H(X|Y,Q). We also use notations
@@
Zipf‘s law. The oldest known of quantitative linguistic laws, Zipf‘s law asserts that, for texts in natural language, the frequency of the $k$-th most common type of word decays approximately as $k^{-\alpha}$, where $\alpha\approx 1$. This regularity was noticed over one hundred years ago [31, 17, 90]. An empirical study of this law across one hundred languages can be found in [63]. Variants of Zipf‘s law with double regimes are also well known [34, 69, 72, 39, 86]. Similar power-law distributions are observed also in ecology, sociology, economics, and physics [91], being a hallmark of complex systems. The literature on Zipf‘s law is scattered over diverse venues. Historical references can be found in [59]. The departure point of a statistically informed theory is [52, 4].
In spite of sheer literature coverage, there is no single explanation of Zipf‘s law. The law can be explained by diverse mechanisms ranging from monkey-typing [62, 68], through preferential attachment [80] and game theory [41], to potential links with semantics and information theory [26]. What matters for our purposes is that stochastic processes $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers with the Zipf probability distribution
@@
provide a compact description of heavy-tailed data distributions. Observation (4) will be our departure point.
@@
Heaps‘ law is widely viewed as a direct consequence of Zipf‘s law with $\beta=1/\alpha$ for $X_{t}=K_{t}$ [62] but the linguistic phenomenology and the statistical theory are more complicated [52, 4, 66, 67, 39, 35, 20, 29]. Understanding precisely what form Zipf‘s law implies a particular form of Heaps‘ law is an important step in our approach.
@@
[…]
@@
A particularly salient empirical finding is that the predictive performance of these models improves with scale in a power-law fashion. The neural scaling law characterizes how the loss function of a foundation model—typically measured by the cross entropy rate on a test dataset—decreases as training data $t$, model size $n$, and compute $c$ increase [51, 44, 46, 48, 58]. Simplifying particular empirical observations and ignoring complex interactions between $t$, $n$, and $c$, this power-law relationship can be approximated as
@@
$h(s,t,n,c):=\frac{\operatorname{\mathbf{E}}(-\log Q_{tnc}(X_{t+1}^{t+s}))}{s}$ (11)
is the expected cross entropy rate of a foundation model $Q_{tnc}$ tested on data $X_{t+1}^{t+s}$ and trained on data $X_{1}^{t}$ with the amount of parameters $n$ and the amount of compute $c$. The exponents satisfy $\gamma_{i}>0$.
Empirically, it is confirmed that $\gamma_{T}>\gamma_{N}$ [51]. This case is called the overparameterization. The underparameterization is the opposite regime $\gamma_{T}<\gamma_{N}$. In this paper, we will derive the neural scaling law from Hilberg‘s hypothesis but our derivation predicts underparameterization if individual parameters have a sandwich bounded Shannon entropy.
@@
We will systematically prove connections among four statistical laws in particular settings that have not been discussed so far. The present paper develops and strengthens the ideas from an earlier and unpublished attempt to attack this topic [28]. We would like to supply a simple-minded baseline that, in contrast to [28], takes into account also the effect of limited amount of training compute $c$. The main contributions of the present paper are organized as a chain of following derivations:
Zipf‘s law $\stackrel{{\scriptstyle(A)}}{{\Rightarrow}}$ Heaps‘ law $\stackrel{{\scriptstyle(B)}}{{\Rightarrow}}$ Hilberg‘s law $\stackrel{{\scriptstyle(C)}}{{\Rightarrow}}$ neural scaling.
Each of the above implications involves different assumptions but we strive at the most general settings such as non-stationary processes. We remark that the conditions for implication $(C)$ (non-stationarity) are more general than that for result $(B)$ (stationarity and Santa Fe decomposition), and those are more general than the requirements for implication $(A)$ (stationarity and mixing). We also notice that the derivation of the neural scaling law rests on the differential Heaps and Hilberg laws (15)–(16), which strengthen the plain Heaps and Hilberg laws (5)–(6). Noticing a practical need for such differential laws seems a novel idea of this paper.
@@
Finally, we show that this form Heaps‘ law is satisfied by sufficiently strongly mixing narrations with sufficiently heavy-tailed marginal distributions. In particular, we derive result $(A)$ for a stationary process $(K_{t})_{t\in\mathbb{Z}}$ with a mixing condition $\Psi<\infty$, where
$\Psi$ $:=\sup_{t,k}\frac{\sum_{i=1}^{t}\sum_{j=1}^{t}P(K_{i}=k,K_{j}=k)-tP(K_{0}=k)}{t(t-1)(P(K_{0}=k))^{2}},$ (17)
and an approximate Zipf law
@@
  • Overparameterization. The result stated in equation (14) suggests theoretical optimality of underparameterization $\gamma_{T}\leq\gamma_{N}$, contrary to the empirical optimality of overparameterization $\gamma_{T}>\gamma_{N}$ [51]. Indeed, suppose that (14) holds. Then we derive
@@
    Thus understanding why overparameterized models perform empirically better requires an explanation. One such explanation may be that condition (13) does not capture the fact that a single real-valued parameter (a single weight) may carry potentially an infinite amount of information (theoretically, it is an infinite sequence of bits). Thus, the effective complexity of the model is greater than the number of real-valued parameters. The empirical advantage of overparameterization may mean that larger models effectively use fewer digits of the binary expansion per real-valued parameter. If this is true then the established notion of overparameterization may be misleading and quantifying the true degrees of freedom remains an open question.
  • Arbitrariness of word shapes. Having a general implication from Hilberg‘s law to neural scaling, the burden of theoretical explanations shifts to stating why Hilberg‘s hypothesis may be sound. In general, Hilberg‘s hypothesis can be viewed as a sort of Zipf‘s law for tokens that are internally random enough like in the Santa Fe process (7). The internal randomness of these tokens can be potentially connected to semantic or linguistic considerations [25, 26, 64]. In particular, we may expect that Hilberg‘s law arises if we have Zipf‘s law for words that have sufficiently arbitrary shapes, cf. [22]. Proving such a claim formally requires a longer excursion to universal coding and involves results that exceed those of papers [25, 26, 27]. For this reason, we postpone this theoretical development to another article.
@@
The organization of this paper is as follows. Section 2 introduces notation and preliminaries. In particular, Section 2.1 treats fundamental inequalities for Shannon entropy and expected cardinality. Section 2.2 discusses the rates of entropy and expected cardinality for arbitrary (non-stationary) processes. Section 2.3 handles the block entropy and the expected block cardinality for stationary processes. Section 2.4 analyzes the vocabulary spectrum elements for IID processes. Section 3 develops the main chain of implications. In particular, Section 3.1 presents the derivation of Heaps‘ law from Zipf‘s law. Section 3.2 develops the implication from Heaps‘ law to Hilberg‘s hypothesis. Section 3.3 establishes the derivation of neural scaling from Hilberg‘s hypothesis. Section 4 concludes the article.
@@ 2.1 Fundamental inequalities
We will be developing parallel results for the Shannon entropy and the expected number of types. Our reasoning is based on simple but systematic information-theoretic considerations and the general spirit of the $I$-measure [49]. For the textbook treatment and the background, we refer to [18, 88]. A handy tool is also a generalization of Shannon information measures to arbitrary $\sigma$-fields [26, Sections 5.3 and 5.4].
@@
$=\operatorname{\mathbf{E}}\left(-\log P(X|Q)\right)$
$=H(X|Q)$
$\geq H(X)-H(Q),$ (30)
@@ 2.3 Stationary processes
Besides the expected number of all types $V(t)$, let us introduce expected spectrum elements $V(t|m)$, defined as the expected number of types that occur exactly $m$ times.
@@
The expected spectrum elements are defined as
$V(t|m):=\sum_{k}P\left(\sum_{i=1}^{t}{\bf 1}{\left\{K_{i}=k\right\}}=m\right).$ (51)
@@ 2.4 IID processes
So far, the statements for Shannon entropy and expected cardinality were mirror-like. However, for more specific processes, the analogy between these two functionals is rather as follows: The expected cardinality applies to IID processes as the Shannon entropy applies to exchangeable Santa Fe processes. This analogy will generalized formally in Section 3.2. Now let us comment on the expected cardinality for IID processes, as it is simpler to analyze. We notice that the expected cardinality for IID sources enjoys further properties: It is a Hausdorff sequence—a discrete-time analog of the Bernstein function.
Combining creatively several sources such as [4, 2, 76], the development is as follows. By the Bernoulli scheme, for an IID process $(K_{t})_{t\in\mathbb{Z}}$, we have
$V(t)$ $=\sum_{k}(1-(1-p_{k})^{t}),$ (54)
$V(t|m)$ $=\begin{cases}\sum_{k}\binom{t}{m}p_{k}^{m}(1-p_{k})^{t-m},&1\leq m\leq t,\\ 0,&m>t,\end{cases}$ (55)
where $p_{k}:=P(K_{t}=k)$. We notice consistency conditions
$\sum_{m=1}^{\infty}V(t|m)$ $=V(t),$ $\sum_{m=1}^{\infty}mV(t|m)$ $=t.$ (56)
Moreover, since $\Delta(1-p_{k})^{t}=-p_{k}(1-p_{k})^{t}$, we also observe that the spectrum elements can be computed from the expected number of types by taking differences,
$V(t|m)$ $=\begin{cases}(-1)^{m+1}\binom{t}{m}\Delta^{m}V(t-m),&1\leq m\leq t,\\ 0,&m>t.\end{cases}$ (57)
Expression (57) reminds of a known concept.
Definition 4 (cf. [2]).
A sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is called a Hausdorff sequence if $v(t)\geq 0$ and $(-1)^{m+1}\Delta^{m}v(t)\geq 0$ for all $m\in\mathbb{N}$ and $n\in\mathbb{N}\cup\left\{0\right\}$. A sequence $u:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is called completely alternating if $u(t)=\Delta v(t)$ for a certain Hausdorff sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$. We call a Hausdorff sequence $v(t)$ standard if $v(0)=0$ and $\Delta v(0)=1$.
The name Hausdorff sequence is non-standard itself. We have coined it by an analogy to the standard term Bernstein function, which is the continuous time-analog, applying derivatives rather than differences [76].
@@
Proposition 11 ([2]).
A sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is a Hausdorff sequence if and only if there exists a unique non-negative measure $F$ on $[0,1]$ such that
$v(t)$ $=tF(\left\{0\right\})+\int_{(0,1)}\frac{1-(1-p)^{t}}{p}dF(p)+F(\left\{1\right\}),$ (58)
$\Delta v(t)$ $=F(\left\{0\right\})+\int_{(0,1)}(1-p)^{t}dF(p).$ (59)
Additionally, we see that a Hausdorff sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is standard if and only if $F(\left\{1\right\})=0$ and the measure $F$ is a probability measure, namely,
$F(\left\{0\right\})+\int_{(0,1)}dF(p)=1.$ (60)
We call measure $F$ the Hausdorff measure.
We notice that the Newton formula $(1-\Delta)^{r}=\sum_{m=0}^{r}\binom{r}{m}(-\Delta)^{m}$ can be rewritten in the following form, which resembles the Taylor expansion,
$v(t-r)$ $=\sum_{m=0}^{r}(-1)^{m}\binom{r}{m}\Delta^{m}v(t-m).$ (61)
By formula (61), consistency conditions analogous to (56) can be verified easily. Namely, a Hausdorff sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ with spectrum elements
$v(t|m)$ $:=\begin{cases}(-1)^{m+1}\binom{t}{m}\Delta^{m}v(t-m),&1\leq m\leq t,\\ 0,&m>t.\end{cases}$ (62)
satisfies conditions $\sum_{m=1}^{t}v(t|m)=v(t)$ and $\sum_{m=1}^{t}mv(t|m)=t$ if and only if it is standard.
To complete this discussion, the expected number of types $V(t)$ is a standard Hausdorff sequence and it has a Hausdorff probability measure supported on discrete points, namely,
$F(A)=\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\in A\right\}}.$ (63)

### v2

@@ 1 Introduction
In the present paper, we aim to provide an actually simple but systematic derivation of this chain of implications that extends earlier results in information theory, probability theory, and quantitative linguistics. We will prove formal theorems about as general stochastic processes as possible rather than experiment with particular empirical data. Our goal is to consolidate knowledge of several scientific disciplines by means of mathematical deduction. We try to avoid a too abstract formalism to make this paper more accessible.
The conceptual trajectory is as follows. Beginning with Zipf‘s law for word frequencies, we derive Heaps‘ law for vocabulary growth. From Heaps‘ law we extract Hilberg‘s hypothesis on the sublinear growth of block entropy. Finally, we show that Hilberg‘s hypothesis leads to the neural scaling that ties model performance to the amounts of training data, model parameters, and training compute. In parallel, we discuss Santa Fe processes—toy stochastic sources that exhibit these statistical laws. Each link in this chain introduces its own assumptions. By isolating these steps, we hope to illuminate where the derivations might be strengthened in the future.
@@
Notation $(X_{t})_{t\in\mathbb{Z}}$ denotes an integer-time countable-alphabet stochastic process, $X_{j}^{k}:=(X_{j},X_{j+1},\ldots,X_{k})$ is a string of random tokens, and $\mathcal{X}_{j}^{k}:=\left\{X_{j},X_{j+1},\ldots,X_{k}\right\}$ is a random set of types. We adopt that $X_{1}^{0}$ is the empty string and $\mathcal{X}_{1}^{0}$ is the empty set. Expectation with respect to probability measure $P$ is denoted $\operatorname{\mathbf{E}}X:=\int XdP$, $H(X):=\operatorname{\mathbf{E}}(-\log P(X))$ is the Shannon entropy, and $V(\mathcal{X}):=\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{X}$ is the expected cardinality. The conditional entropy is H(X|Q):=H(X,Q)H(Q)H(X|Q):=H(X,Q)-H(Q), the mutual information is I(X;Y):=H(X)H(X|Y)I(X;Y):=H(X)-H(X|Y), and the conditional mutual information is I(X;Y|Q):=H(X|Q)H(X|Y,Q)I(X;Y|Q):=H(X|Q)-H(X|Y,Q). We also use
@@
Zipf‘s law. The oldest known of quantitative linguistic laws, Zipf‘s law asserts that, for texts in natural language, the frequency of the $k$-th most common type of word decays approximately as $k^{-\alpha}$, where $\alpha\approx 1$. This regularity was noticed over one century ago [33, 19, 98]. An empirical study of this law across one hundred languages can be found in [67]. Variants of Zipf‘s law with double regimes are also well known [36, 73, 77, 41, 94]. Similar power-law distributions are observed also in ecology, sociology, economics, and physics [99], being a hallmark of complex systems. The literature on Zipf‘s law is scattered over diverse venues. Historical references can be found in [63]. The departure point of a statistically informed theory of Zipf‘s law is formed by references [54, 55, 78, 5].
In spite of sheer literature coverage (or maybe because of that), there is no single explanation of Zipf‘s law. The law can be explained by diverse mechanisms ranging from monkey-typing [66, 72], through preferential attachment [87], also known as the Chinese restaurant process [3, page 92], to potential links with game theory [43], semantics, and information theory [28]. What matters for our purposes is that stochastic processes $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers with the approximate Zipf probability distribution
@@
provide a compact description of heavy-tailed data distributions. Statistical law (4) will be our departure point.
@@
Heaps‘ law is widely viewed as a direct consequence of Zipf‘s law with $\beta=1/\alpha$ for $X_{t}=K_{t}$ [66] but the linguistic phenomenology and the statistical theory are more complicated [54, 55, 5, 70, 71, 41, 37, 22, 16, 31]. Heaps‘ law has been also studied for large language models [90, 58]. Understanding precisely what form of Zipf‘s law implies a particular form of Heaps‘ law is an important step in our approach.
@@
Contrary to intuitions about complex structures, Hilberg‘s law arises also in highly simplified settings of Santa Fe processes. In particular, it suffices to assume that narration $(K_{t})_{t\in\mathbb{Z}}$ is an IID source with the Zipf distribution (4) and knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent fair coin flips, independent of narration $(K_{t})_{t\in\mathbb{Z}}$. Under these conditions, we obtain Heaps‘ law (5) and Hilberg‘s law (6) with $\beta=1/\alpha$ [27, 28].
@@
A particularly salient empirical finding is that the predictive performance of these models improves with scale in a power-law fashion. The neural scaling law characterizes how the loss function of a foundation model—typically measured by the cross entropy rate on a test dataset—decreases as training data $t$, model size $n$, and compute $c$ increase [53, 46, 48, 50, 76, 62]. Simplifying particular empirical observations and ignoring complex interactions between $t$, $n$, and $c$, this power-law relationship can be approximated as
@@
$h(s,t,n,c):=\sup_{k\geq t}\frac{\operatorname{\mathbf{E}}(-\log Q_{tnc}(X_{k+1}^{k+s}))}{s}$ (11)
is the worst-case expected cross entropy rate of a foundation model $Q_{tnc}$ tested on data $X_{k+1}^{k+s}$ and trained on data $X_{1}^{t}$ with the amount of parameters $n$ and the amount of compute $c$. The exponents satisfy $\gamma_{i}>0$.
Empirically, it is confirmed that $\gamma_{T}>\gamma_{N}$ [53, 50, 76]. This case is called the overparameterization. The underparameterization is the opposite regime $\gamma_{T}<\gamma_{N}$. In this paper, we will derive the neural scaling law from Hilberg‘s hypothesis but our derivation predicts underparameterization if individual parameters have a sandwich bounded Shannon entropy.
@@
We will systematically prove connections among four statistical laws in particular settings that have not been discussed so far. The present paper develops and strengthens the ideas from an earlier unpublished attempt to attack this topic [30]. We would like to supply a simple-minded baseline that, in contrast to [30], takes into account also the effect of limited amount of training compute $c$. The main contributions of the present paper are organized as a chain of following derivations:
Zipf‘s law $\stackrel{{\scriptstyle(A)}}{{\implies}}$ Heaps‘ law $\stackrel{{\scriptstyle(B)}}{{\implies}}$ Hilberg‘s law $\stackrel{{\scriptstyle(C)}}{{\implies}}$ neural scaling.
Each of the above implications involves different assumptions but we strive at the most general settings—such as non-stationary processes. In particular, the conditions for implication $(C)$ (arbitrary non-stationarity) are more general than that for result $(B)$ (stationarity and Santa Fe decomposition), and those are more general than the requirements for implication $(A)$ (stationarity and mixing). We observe that the derivation of neural scaling rests on the differential Heaps and Hilberg laws (15)–(16), which strengthen the plain Heaps and Hilberg laws (5)–(6). Noticing a practical need for such differential laws seems a novel idea of this paper. It would be nice to combine this idea with the powerful IID framework by Karlin [54] in the future, possibly also extending it from IID to mixing or ergodic processes.
@@
Finally, we show that this form of Heaps‘ law is satisfied by sufficiently strongly mixing narrations with sufficiently heavy-tailed marginal distributions. In particular, we derive result $(A)$ for a stationary process $(K_{t})_{t\in\mathbb{Z}}$ such that
$\sum_{k\in\mathbb{N}}P(K_{0}=k){\bf 1}{\left\{P(K_{0}=k)\leq p\right\}}$ $\stackrel{{\scriptstyle*}}{{>}}p^{1-\beta},$ (17)
$\frac{P(K_{t}=k|K_{0}=k)}{P(K_{0}=k)}$ $\stackrel{{\scriptstyle*}}{{<}}1.$ (18)
In particular, condition (17) holds for an approximate Zipf law
@@
Condition (18) is satisfied for IID and finite-state processes. It is implied by condition $\psi^{*}(1)<\infty$ in the terminology of Bradley [13]. We also note that
$\lim_{t\to\infty}\frac{P(K_{t}=k|K_{0}=k)}{P(K_{0}=k)}=1$ (20)
holds for any stationary strongly mixing process $(K_{t})_{t\in\mathbb{Z}}$. For a comprehensive survey of various mixing conditions, we refer to [13].
@@
  • Tightness of the derived bounds. Suppose that the neural scaling law of form (14) holds. Then we derive inequalities
@@
    In particular, for the PPM-based estimate $\beta\approx 0.8$ reported by Takahira et al. [88], we obtain $\gamma_{T}\leq 0.2$ and $\gamma_{N}\leq 0.25$. This is quite a loose upper bound of the values $\gamma_{T}\approx 0.095$ and $\gamma_{N}\approx 0.076$ reported by Kaplan et al. [53]. We note that these estimates were obtained for corpora of a different magnitude—gigabytes of tokens in the case of [88] to be contrasted with terabytes of tokens in the case of [53]. Further data analysis is needed to explain the large gap between the empirical values of exponents $\gamma_{T}$ and $\gamma_{N}$ and their upper bounds stemming from known estimates of the Hilberg exponent $\beta$. It is plausible that $\beta$ is substantially larger than $0.8$ for internet-sized corpora.
  • Overparameterization. The derived inequalities (21)–(22) suggest theoretical optimality of underparameterization $\gamma_{T}\leq\gamma_{N}$, contrary to the empirical optimality of overparameterization $\gamma_{T}>\gamma_{N}$ [53, 50, 76]. Thus understanding why overparameterized models perform empirically better may require an explanation. One such explanation may be that condition (13) does not capture the fact that a single real-valued parameter (a single weight) may carry potentially an infinite amount of information (theoretically, it is an infinite sequence of bits). Thus, the effective complexity of the model is greater than the number of real-valued parameters. The empirical advantage of overparameterization may mean that larger models effectively use fewer digits of the binary expansion per real-valued parameter. If this is true then the established notion of overparameterization may be misleading and quantifying the true degrees of freedom remains an open question.
  • Arbitrariness of word shapes. Having a general implication from Hilberg‘s law to neural scaling, the burden of theoretical explanations shifts to stating why Hilberg‘s hypothesis may be sound. In general, Hilberg‘s hypothesis can be viewed as a sort of Zipf‘s law for tokens that are internally random enough like in the Santa Fe process (7). The internal randomness of these tokens can be potentially connected to semantic or linguistic considerations [27, 28, 68]. In particular, we may expect that Hilberg‘s law is equivalent to Zipf‘s law for words that have sufficiently arbitrary shapes. Arbitrariness of word shapes is one of the classical tenets of linguistics [24]. Proving this equivalence formally requires a longer excursion to universal coding and involves results that exceed those of papers [27, 28, 29]. For this reason, we postpone this theoretical development to another article.
@@
The organization of this paper is as follows. Section 2 introduces notation and preliminaries. In particular, Section 2.1 treats fundamental inequalities for Shannon entropy and expected cardinality. Section 2.2 discusses the rates of entropy and expected cardinality for arbitrary (non-stationary) processes. Section 2.3 handles the block entropy and the expected block cardinality for stationary processes. It also introduces spectrum elements and bounds the rate of hapaxes. Section 2.4 analyzes the spectrum elements for IID processes. Section 3 develops the main chain of implications. In particular, Section 3.1 presents the derivation of Heaps‘ law from Zipf‘s law. Section 3.2 develops the implication from Heaps‘ law to Hilberg‘s hypothesis. Section 3.3 establishes the derivation of neural scaling from Hilberg‘s hypothesis. Section 4 concludes the article.
@@ 2.1 Fundamental inequalities
We will be developing parallel results for the Shannon entropy and the expected number of types. Our reasoning is based on simple but systematic information-theoretic considerations and the general spirit of the $I$-measure [51]. For the textbook treatment and the background, we refer to [20, 96]. A handy tool is also a generalization of Shannon information measures to arbitrary $\sigma$-fields [28, Sections 5.3 and 5.4].
@@
$=\operatorname{\mathbf{E}}\left(-\log P(X|Q)\right)=H(X|Q),$ (32)
@@ 2.3 Stationary processes
Besides the expected number of all types $V(t)$, let us introduce spectrum elements $V(t|m)$, defined as the expected number of types that occur exactly $m$ times [54, 55, 5].
@@
The spectrum elements are defined as
$V(t|m)$ $:=\sum_{k\in\mathbb{N}}P\left(\sum_{i=1}^{t}{\bf 1}{\left\{K_{i}=k\right\}}=m\right),$ $1$ $\leq m\leq t.$ (53)
@@ 2.4 IID processes
So far, the statements for Shannon entropy and expected cardinality were mirror-like. However, for more specific processes, the analogy between these two functionals is rather as follows: The expected cardinality applies to IID processes in a similar fashion as the Shannon entropy applies to exchangeable Santa Fe processes. This analogy will be generalized in Section 3.2. Now we consider the expected cardinality for IID processes, being simpler to analyze. We notice that the expected cardinality for IID sources enjoys further properties: It is a Hausdorff sequence—a discrete-time analog of a Bernstein function. This observation nicely complements the results by Karlin [54] for IID and Poisson cases.
The development is as follows.
Definition 4.
A sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is called a Hausdorff sequence if $v(t)\geq 0$ and $(-1)^{m+1}\Delta^{m}v(t)\geq 0$ for all $m\in\mathbb{N}$ and $n\in\mathbb{N}\cup\left\{0\right\}$. A sequence $u:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is called completely alternating if $u(t)=\Delta v(t)$ for a certain Hausdorff sequence $v(t)$. We call a sequence $v(t)$ standard if $v(0)=0$ and $\Delta v(0)=1$.
Remark: The name Hausdorff sequence is non-standard itself. We have coined it by an analogy to the standard term Bernstein function, which is the continuous time-analog, applying derivatives rather than differences [83].
@@
Proposition 11.
A sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$ is a Hausdorff sequence if and only if there exists a unique non-negative measure $\tilde{v}$ on $[0,1]$ such that
$v(t)$ $=t\tilde{v}(\left\{0\right\})+\int_{(0,1)}\frac{1-(1-p)^{t}}{p}d\tilde{v}(p)+\tilde{v}(\left\{1\right\}),$ (56)
$\Delta v(t)$ $=\tilde{v}(\left\{0\right\})+\int_{(0,1)}(1-p)^{t}d\tilde{v}(p).$ (57)
Remark: We call measure $\tilde{v}$ the Hausdorff measure of sequence $v(t)$. A Hausdorff sequence $v(t)$ is standard if and only if $\tilde{v}(\left\{1\right\})=0$ and the measure $\tilde{v}$ is a probability measure.
Proof.
See [44, 2]. ∎
Definition 5.
For an arbitrary sequence $v:\mathbb{N}\cup\left\{0\right\}\to\mathbb{R}$, we define its Taylor elements
$v(t\|m)$ $:=(-1)^{m+1}\binom{t}{m}\Delta^{m}v(t-m),$ $1$ $\leq m\leq t.$ (58)
We notice that $v(t\|m)\geq 0$ if and only if sequence $v(t)$ is a Hausdorff sequence.
Proposition 12.
The Taylor elements satisfy consistency conditions
$\sum_{m=1}^{\infty}v(t\|m)$ $=v(t),$ $\sum_{m=1}^{\infty}mv(t\|m)$ $=t.$ (59)
if and only if sequence $v(t)$ is standard.
Proof.
The claim follows by the Newton formula $(1-\Delta)^{r}=\sum_{m=0}^{r}\binom{r}{m}(-\Delta)^{m}$, written as
$v(t-r)$ $=\sum_{m=0}^{r}(-1)^{m}\binom{r}{m}\Delta^{m}v(t-m),$ (60)
resembling the Taylor expansion. It suffices to consider (60) for $r=t$ and r=t1r=t-1. ∎
Now, without loss of generality, let us assume that the alphabet of an IID process $(K_{t})_{t\in\mathbb{Z}}$ is the set of natural numbers. By the Bernoulli scheme, the expected number of types and the spectrum elements defined via (53) equal
$V(t)$ $=\sum_{k\in\mathbb{N}}(1-(1-p_{k})^{t}),$ (61)
$V(t|m)$ $=\sum_{k\in\mathbb{N}}\binom{t}{m}p_{k}^{m}(1-p_{k})^{t-m},$ $1$ $\leq m\leq t,$ (62)
where $p_{k}:=P(K_{0}=k)$. By formula (62) and identity $\Delta(1-p_{k})^{t}=-p_{k}(1-p_{k})^{t}$, the spectrum elements $V(t|m)$ and the Taylor elements of $V(t)$ are equal,
$V(t|m)=V(t\|m).$ (63)
Consequently, sequence $V(t)$ is a Hausdorff sequence. Moreover, sequence $V(t)$ is standard and spectrum elements $V(t|m)$ satisfy consistency conditions analogous to (59). The Hausdorff measure of $V(t)$ is an atomic probability measure and assumes form
$\tilde{V}(A)=\sum_{k\in\mathbb{N}}p_{k}{\bf 1}{\left\{p_{k}\in A\right\}}.$ (64)